Q. Two fair dice are thrown together. What is the probability that at least one die shows 6?

  • (A) 11/36
  • (B) 25/36
  • (C) 1/6
  • (D) 5/36
βœ… Correct Answer: (A) 11/36
Explanation: Using complement: P(no 6) = (5/6)^2 = 25/36. Required probability = 1 - 25/36 = 11/36.
Explanation by: Krishna Sharma
Using complement: P(no 6) = (5/6)^2 = 25/36. Required probability = 1 - 25/36 = 11/36.

πŸ“ Notes by Other Students/ Teachers


πŸ“Š Question Analytics

πŸ‘οΈ
43
Total Visits
πŸ“½οΈ
4 mo ago
Published
πŸŽ–οΈ
Krishna Sharma
Publisher
πŸ“ˆ
91%
Success Rate